$SPAD/src/input f04adf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)cl all
showArrayValues true
showScalarValues true
ia:=3
b:Matrix Complex SF:=
 [[1 ,0 ,0 ]]
ib:=3
n:=3
m:=1
ic:=3
a:Matrix Complex SF:=
 [[1,1 + 2*%i,2 + 10*%i],_
  [1 + %i,3*%i,-5 + 14*%i],_
  [1 + %i,5*%i,-8 + 20*%i]]
result:=f04adf(ia,b,ib,n,m,ic,a,-1)

———-

References^{[1] nothing}

4